# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

### v3

@@ 4 Extended $\mathcal{K}\mathcal{L}$ Theory for Discrete-Time Systems Maximization
In this section, we come back to the initial problem that is a peaks computation problem i.e. the maximization of a function over the reachable values of a discrete-time dynamical system. We recall that the data are the following:
@@
We recall that we want compute $\nu_{\rm opt}=\sup\{\nu_{k}:k\in\mathbb{N}\}$ where $\nu_{k}=\sup\{\varphi(T^{k}(x):x\in X^{\mathrm{in}}\}$. To use the results of Section 3 with the sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$, we suppose that Assumption 1 holds.
We propose to simplify the notation of the maximal value of an objective function $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ on $X^{\mathrm{in}}$, we introduce the following new notation:
@@
and the set of functions which are bounded from above on $X^{\mathrm{in}}$:
$\operatorname{Fin}\left(X^{\mathrm{in}}\right):=\{f:\mathbb{R}^{d}\mapsto\mathbb{R}\cup\{+\infty\}:\overline{f_{X^{\mathrm{in}}}}<+\infty\}$
With these notations, $\nu_{k}=\overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}$ and by following Assumption 1, we have for all $k\in\mathbb{N}$, $\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$. Following Subsection 3.2, we have to study, first, the limsup of the sequence $\nu$ and then construct, when the limsup is finite, an element $(h,\beta)$ of $\mathbb{\Gamma}(\nu)$ i.e. $h$ is strictly increasing and continuous on $[0,1]$, $\beta\in(0,1)$ and for all $k\in\mathbb{N}$, $h(\beta^{k})$ is greater than $\nu_{k}$. We also have to guarantee that if ${\mathrm{R}}_{\nu}^{s}$ is nonempty, $h$ has to be useful for $\nu$ i.e. $\mathcal{S}(\nu,h)$ is nonempty. Without this latter condition, the function $h$ is so useless as $\mathfrak{F}_{u}(k,h,\beta)=+\infty$. In this section, we will revisit $\mathcal{KL}$ class of functions to study the limsup of $\nu$ and to construct our necessary pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$.
4.1 Using the classical $\mathcal{KL}$ class of functions
To study global asymptotic stability of discrete-time systems can be done by means of comparison functions (see e.g. [GGLW14]). For a comprehensive literature about comparison functions, the interested reader can refer to  [Kel14]. Comparison functions can be used for our problem to study the limsup of the sequence $\nu$.
We start by recalling some necessary definitions about comparison functions.
Definition 2 ($\mathcal{K}$/$\mathcal{K}_{\infty}$ functions classes).
A function $\alpha:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ belongs to $\mathcal{K}$ if $\alpha$ is strictly increasing continuous and $\alpha(0)=0$. If moreover, $\lim_{x\to+\infty}\alpha(x)=+\infty$, $\alpha$ is said to belong to $\mathcal{K}_{\infty}$.
Definition 3 ($\mathcal{L}$ functions class).
A function $\sigma:\mathbb{R}_{+}\to\mathbb{R}_{+}$ belongs to $\mathcal{L}$ if it is continuous, strictly decreasing and $\lim_{x\to+\infty}\sigma(x)=0$.
Definition 4 ($\mathcal{K}\mathcal{L}$ functions class).
A function $\gamma:\mathbb{R}_{+}\times\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ belongs to $\mathcal{K}\mathcal{L}$ if for all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)$ belongs to $\mathcal{K}$ and if for all $s\in\mathbb{R}_{+}$, $t\mapsto\gamma(s,t)$ belongs to $\mathcal{L}$.
The class $\mathcal{K}$ is included in $\Omega([0,1])$ but $\mathcal{K}$ functions are strictly positive except at 0 where they are null. Hence, $h\in\mathcal{K}\cap\mathbb{\Gamma}_{\rm fun}(\nu)$ is useful for $\nu$ if at least one term $\nu_{k}$ is strictly positive. Thus, in this subsection, Assumption 2 must hold.
Assumption 2.
There exists $k\in\mathbb{N}$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{k}(x))>0$.
We recall the result obtained by Sontag.
Lemma 1 (Proposition 7 of [Son98]).
For all $\gamma\in\mathcal{K}\mathcal{L}$, there exists $\theta_{1},\theta_{2}\in\mathcal{K}_{\infty}$ such that for all $s,t\in\mathbb{R}_{+}$, $\gamma(s,t)\leq\theta_{1}(\theta_{2}(s)e^{-t})$.
Theorem 4.
Assume that there exist a function $\gamma\in\mathcal{K}\mathcal{L}$ and a function $\psi:\mathbb{R}^{d}\mapsto\mathbb{R}_{+}$ such that $\overline{\psi_{X^{\mathrm{in}}}}>0$. We also assume that for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$:
$\varphi(T^{k}(x))\leq\gamma(\psi(x),k)\enspace.$
Then:
  1. 1.
    $\limsup_{n\to+\infty}\nu_{n}\leq 0$ and if, moreover, Assumption 2 holds then ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$;
  2. 2.
    there exists $h\in\mathcal{K}_{\infty}$ s.t. $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$ and, if, moreover, Assumption 2 holds then $h$ is useful for $\nu$.
Proof.
[…]
2. This a direct consequence of the Sontag’s result (Lemma 1 here). Indeed, there exist $\theta_{1},\theta_{2}\in\mathcal{K}_{\infty}$ such that for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$ $\varphi(T^{k}(x))\leq\gamma(\psi(x),k)\leq\theta_{1}(\theta_{2}(\psi(x))e^{-k})$. Then $\nu_{k}\leq h(e^{-1}):=\sup_{x\in X^{\mathrm{in}}}\theta_{1}(\theta_{2}(\psi(x)e^{-k}))$. It is easy to see, as $\theta_{1}$ and $\theta_{2}$ are increasing and continuous, that $h$ can be also written $y\mapsto\theta_{1}(\theta_{2}(\overline{\psi_{X^{\mathrm{in}}}}y))$. The function $h$ is thus the compositions of $\theta_{1}$, $\theta_{2}$ and $y\mapsto\overline{\psi_{X^{\mathrm{in}}}}y$ which belong to $\mathcal{K}_{\infty}$ (the third as $\overline{\psi_{X^{\mathrm{in}}}}>0$) and thus belongs to $\mathcal{K}_{\infty}$. Finally, under Assumption 2, there exists $k\in\mathbb{N}$, such that $\nu_{k}>0=h(0)$. ∎
Usually i.e. when $\varphi=\|\cdot\|$, the inequality $\varphi(T^{k}(x))\leq\gamma(\psi(x),k)$ (for $\psi=\|\cdot\|$) is true for all $x\in\mathbb{R}^{d}$ and all $k\in\mathbb{N}$ if and only if the system $x_{k+1}=T(x_{k}),x_{0}\in\mathbb{R}^{d}$ is said to be uniformly global asymptotic stability (UGAS) (see for example [NT04, Prop. 1] or [ATP10, Th. 2]. This implies that $x_{k}$ tends to 0 as $k$ tends to +$\infty$+ ∞. Thus, with a general function $\varphi$, we obtain a one-sided information on the asymptotic behaviour of the dynamics. We could obtain the other side applying the theorem for $-\varphi$.
@@ 4.2 Revisiting $\mathcal{KL}$ functions : the class $\mathcal{KL}_{\rm gen}$
In this paper, we are not interested in stability of the dynamical system, we can relax the strict increase (or decrease), the continuity and the positivity (on the state variable) conditions required for $\mathcal{KL}$ functions.
@@
A function $\gamma:\mathbb{R}\times\mathbb{R}_{+}\mapsto\mathbb{R}$ belongs to $\mathcal{KL}_{\rm gen}$ if for all $s\in\mathbb{R}$, $t\mapsto\gamma(s,t)$ is increasing and for all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)$ is decreasing.
@@
A pair $(\gamma,\theta)\in\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ if and only if :
@@
Moreover, $(\gamma,\theta)$ is said to useful for $\nu$ such that ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ if and only if the set:
@@
Following the notations introduced earlier, we have the equivalence:
$\nu\in\Lambda\iff\sup_{k\in\mathbb{N}}\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$
From Prop. 1 we have the equivalence, we also have:
$\nu\in\Lambda\iff\limsup_{k\to+\infty}\nu_{k}=\limsup_{k\to+\infty}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))=\limsup_{k\to+\infty}\ \overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}\in\mathbb{R}\cup\{-\infty\}$
We can establish a first result which states that the existence of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds is equivalent to the fact that the sequence $\nu$ is bounded from above. This is an adaptation of classical results dealing with stability and $\mathcal{KL}$ upper bounds.
@@
The sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$ belongs to $\Lambda$ if and only if there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. If, moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ then $(\gamma,\theta)$ can be chosen to be useful for $\nu$.

### v4

@@ 4 Extended $\mathcal{K}\mathcal{L}$ Theory for Discrete-Time Systems Maximization
In this paper, we are interested in the resolution of Problem (DPCP), for which the underlying dynamical system $x_{k+1}=T(x_{k})$ is not necessarily asymptotically stable, as shown in the example provided in Section 3. However, Inequality (6) holds if and only if the system $x_{k+1}=T(x_{k}),x_{0}\in\mathbb{R}^{d}$ is said to be uniformly global asymptotic stability (UGAS) (see, for example [NT04, Prop. 1] or [ATP10, Th. 2]). On the other hand, Inequality (10) provides useful information on the asymptotic behavior of the system: the limit superior of $\nu_{k}$ is smaller than $\inf_{k\in\mathbb{N}}\gamma(\sup_{x\in X^{\mathrm{in}}}\psi(x),k)$ and thus finite. However, Inequality (10) forces our sequence $\nu$ to have a positive term to obtain the usefulness of the pair proposed in (11). Therefore, Inequality (10) constitutes our starting point. Thus, we define the class of functions $\mathcal{KL}_{\rm gen}$ that extends the class of functions $\mathcal{K}\mathcal{L}$. The class $\mathcal{KL}_{\rm gen}$ permits the construction of a useful element $(h,\beta)\in\mathbb{\Gamma}(\nu)$ when ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ and the analog of Inequality (10) is verified by replacing $\gamma$ with a $\mathcal{KL}_{\rm gen}$ function. Contrary to the classical $\mathcal{K}\mathcal{L}$ certificate approach, we prove a converse theorem: if ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, then the analog of Inequality (10) holds for some $\gamma\in\mathcal{KL}_{\rm gen}$. We also provide an extension to the famous Sontag’s result [Son98, Proposition 7] recalled above for the $\mathcal{KL}_{\rm gen}$ class. It allows to obtain an element $h\in\mathbb{\Gamma}(\nu)$ such that $e^{-1}\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$ when the analogue of (10) is satisfied.
We recall that the data are as follows:
@@
We recall that we define $\nu_{\rm opt}=\sup\{\nu_{k}:k\in\mathbb{N}\}$ where $\nu_{k}=\sup\{\varphi(T^{k}(x):x\in X^{\mathrm{in}}\}$. Naturally, we suppose that Assumption 2 holds. To simplify the notation of the maximal value on $X^{\mathrm{in}}$ of a function $f:\mathbb{R}^{d}\mapsto\mathbb{R}\cup\{+\infty\}$, we introduce the following new notation:
@@
and the set of functions that are bounded from above on $X^{\mathrm{in}}$:
$\operatorname{Fin}\left(X^{\mathrm{in}}\right):=\{f:\mathbb{R}^{d}\mapsto\mathbb{R}\cup\{+\infty\}:\overline{f_{X^{\mathrm{in}}}}<+\infty\}.$
With those notations, $\nu_{k}=\overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}$ and by following Assumption 2, we have for all $k\in\mathbb{N}$, $\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$.
@@ 4.1 Revisiting $\mathcal{KL}$ functions : the class $\mathcal{KL}_{\rm gen}$
One of the limitations of using classical $\mathcal{K}\mathcal{L}$ functions in (10) is the positivity of $\mathcal{K}\mathcal{L}$ functions. To ensure the usefulness of the element $(h,\beta)$ of Equation (11), our sequence $\nu$ must have a positive term. Then, we should relax the positivity on the state-variable part of $\mathcal{K}\mathcal{L}$ functions (and not on the time variable). We also relax the strict increase (or decrease) and continuity required to be a $\mathcal{KL}$ function.
@@
A function $\gamma:\mathbb{R}\times\mathbb{R}_{+}\mapsto\mathbb{R}$ belongs to $\mathcal{KL}_{\rm gen}$ if
  • for all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)$ is increasing;
  • for all $s\in\mathbb{R}$, $t\mapsto\gamma(s,t)$ is decreasing.
Example 1 (A pertinent $\mathcal{KL}_{\rm gen}$ for Leslie models).
For the example introduced in Section 3, we compute a $\mathcal{KL}_{\rm gen}$ upper bound for $(X_{\ell},L,\varphi_{\ell})$. In this example, we construct a family of $\mathcal{KL}_{\rm gen}$ functions for which some of them will be $\mathcal{KL}_{\rm gen}$ upper bounds for $(X_{\ell},L,\varphi_{\ell})$. For this construction, we introduce the following auxiliary function.
$H:\{(x,y)\in\mathbb{R}^{2}:x+y\neq 0\}\ni(u,v)\mapsto\dfrac{\lambda_{2}}{c}\dfrac{\dfrac{\lambda_{1}}{\lambda_{2}}v+u}{v+u}.$ (26)
We observe that we have for all $x\in X_{\ell}$ (defined in (24)) and all $k\in\mathbb{N}$
$\varphi_{\ell}(L^{k}x)=\dfrac{\lambda_{1}^{k+1}\mathfrak{a}(x)+\lambda_{2}^{k+1}\mathfrak{b}(x)}{c\left(\lambda_{1}^{k}\mathfrak{a}(x)+\lambda_{2}^{k}\mathfrak{b}(x)\right)}=\dfrac{\lambda_{2}}{c}\dfrac{\dfrac{\lambda_{1}}{\lambda_{2}}\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}+\dfrac{\mathfrak{b}(x)}{\mathfrak{a}(x)}}{\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}+\dfrac{\mathfrak{b}(x)}{\mathfrak{a}(x)}}=H\left(\dfrac{\mathfrak{b}(x)}{\mathfrak{a}(x)},\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}\right).$ (27)
From (27), we propose to construct a $\mathcal{KL}_{\rm gen}$ function $\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\gamma(s,t)$ from $H$. Indeed, the final purpose of the constructed $\mathcal{KL}_{\rm gen}$ function is to provide an upper bound of $\varphi_{\ell}\circ L^{k}$ on $X_{\ell}$. From the analysis of the sign of the partial derivatives of $H$, we remark that
  • for all $v>0$ (resp. $v<0$), $u\mapsto H(u,v)$ strictly increases on $(-\infty,-v)$ and on $(-v,+\infty)$ (resp. strictly decreases on $(-\infty,-v)$ and on $(-v,+\infty)$);
  • for all $u<0$ (resp. $u>0$), $v\mapsto H(u,v)$ strictly increases on $(-\infty,-u)$ and on $(-u,+\infty)$ (resp. strictly decreases on $(-\infty,-u)$ and on $(-u,+\infty)$).
From these monotonicity properties, $H$ is a candidate to be a $\mathcal{KL}_{\rm gen}$ function on negative values of the state-variable $s$ and with positive second arguments. Therefore, we will consider $|\lambda_{1}/\lambda_{2}|^{t}$ rather than $(\lambda_{1}/\lambda_{2})^{t}$ as second argument in $H$. For positive values $s$, the $\mathcal{KL}_{\rm gen}$ function $\gamma(s,t)$ cannot be equal to $H(s,|\lambda_{1}/\lambda_{2}|^{t})$ and we will propose a strictly increasing affine function. Moreover, $H(s,|\lambda_{1}/\lambda_{2}|^{t})$ is not defined when $s=-|\lambda_{1}/\lambda_{2}|^{t}$ and we have to avoid those points. Finally, the main objective of this $\mathcal{KL}_{\rm gen}$ function is to provide an upper bound over $\varphi_{\ell}\circ L^{k}$ on $X_{\ell}$. Unfortunately, for all $s\in(-|\lambda_{1}/\lambda_{2}|^{t},0)$, $H(s,|\lambda_{1}/\lambda_{2}|^{t})$ is strictly negative and thus does not provide an upper bound for $\varphi_{\ell}\circ L^{k}$ which is strictly positive on $X_{\ell}$. We also remark that $s\mapsto H(s,|\lambda_{1}/\lambda_{2}|^{t})$ is not bounded from above on $(-\infty,-|\lambda_{1}/\lambda_{2}|^{t})$. Consequently, we artificially bound it, allowing only values $s$ smaller than $|\lambda_{1}/\lambda_{2}|^{t}-\varepsilon$ for a given strictly positive $\varepsilon$. We will put constraints over $\varepsilon$ when we impose the $\mathcal{KL}_{\rm gen}$ function to be an upper bound of $\varphi_{\ell}\circ L^{k}$ on $X_{\ell}$. Finally, given $\varepsilon>0$, $H$ is exploited at $(s,|\lambda_{1}/\lambda_{2}|^{t})$ only when $s<-|\lambda_{1}/\lambda_{2}|^{t}-\varepsilon$. Otherwise, we use an affine function increasing in $s$ and in $|\lambda_{1}/\lambda_{2}|^{t}$ proved to be greater than $H(-|\lambda_{1}/\lambda_{2}|^{t}-\varepsilon,|\lambda_{1}/\lambda_{2}|^{t})$. Finally, we get a family of $\mathcal{KL}_{\rm gen}$ functions defined for all $\varepsilon>0$ as follows,
$\gamma_{\varepsilon}:\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\left\{\begin{array}[]{lr}H\left(s,\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}\right)&\text{ if }s\leq-\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}-\varepsilon\\ \left(s+\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}+\varepsilon\right)+H\left(-\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}-\varepsilon,\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}\right)&\text{ otherwise }\end{array}\right$. (28)
We prove in the Appendix that for all $\varepsilon>0$, $\gamma_{\varepsilon}\in\mathcal{KL}_{\rm gen}$.
Note that we could replace $H(-|\lambda_{1}/\lambda_{2}|^{t}-\varepsilon,|\lambda_{1}/\lambda_{2}|^{t})$ by a strictly positive function increasing with respect to $|\lambda_{1}/\lambda_{2}|^{t}$ and greater than $H(-|\lambda_{1}/\lambda_{2}|^{t}-\varepsilon,|\lambda_{1}/\lambda_{2}|^{t}$. The current choice of affine function makes $\gamma_{\varepsilon}$ continuous, which is not mandatory to belong to $\mathcal{KL}_{\rm gen}$.
Applying  (28) to the numerical data of Subsection 3.4, we have
$\gamma_{\varepsilon}:\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\left\{\begin{array}[]{lr}4\dfrac{0.5^{t+1}+s}{0.5^{t}+s}&\text{ if }s\leq-(0.5)^{t}-\varepsilon\\ \\ s+0.5^{t}(1+6\varepsilon^{-1})+\varepsilon+4&\text{ if }s>-0.5^{t}-\varepsilon\end{array}\right$.
@@
A pair $(\gamma,\theta)\in\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ if and only if
@@
Moreover, $(\gamma,\theta)$ is said to be useful for $\nu$ such that ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ if and only if the set
@@
Example 2 (A useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ upper bound for $(X_{\ell},L,\varphi_{\ell})$).
We now propose, as ${\mathrm{R}}_{\omega}^{s}\neq\emptyset$, a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ upper bound for the peak computation problem introduced in ($DPCP_{\ell}$). We show that for well-chosen $\varepsilon$, $\gamma_{\varepsilon}$ introduced in (28) combined with $\theta_{\ell}:=-|\mathfrak{b}/\mathfrak{a}|$ that belongs to $\operatorname{Fin}\left(X_{\ell}\right)$ provide a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ upper bound for $(X_{\ell},L,\varphi_{\ell})$.
The choice of $\varepsilon$ depends on supremum of $\theta_{\ell}$ over $\mathfrak{a}^{-1}((-\infty,0))$ and over $\mathfrak{a}^{-1}((0,+\infty))$. Using the partial derivatives of $|\mathfrak{b}/\mathfrak{a}|$, we conclude that
[…]
For all $x\in X_{\ell}$ such that $\mathfrak{a}(x)<0$, we have $\theta_{\ell}(x)<-1$. Moreover, the set $\mathfrak{a}^{-1}((-\infty,0))$ is nonempty if and only if $\mathfrak{a}(x^{ul})<0$. Thus, the supremum of $\theta_{\ell}$ over $\mathfrak{a}^{-1}((-\infty,0))$ is strictly smaller than -1. For all $x\in X_{\ell}$ such that $0<\mathfrak{a}(x)<\mathfrak{b}(x)$, we also have $\theta_{\ell}(x)<-1$. However, if the set $\{x\in X_{\ell}:\mathfrak{a}(x)\geq\mathfrak{b}(x)\}$ (included in $\mathfrak{a}^{-1}((0,+\infty))$) is nonempty, for all $\varepsilon>0$, we can find $x\in X_{\ell}$ such that $0<\mathfrak{a}(x)<\mathfrak{b}(x)$ and $1-\varepsilon\leq\theta_{\ell}(x)$. Note that $\{x\in X_{\ell}:\mathfrak{a}(x)\geq\mathfrak{b}(x)\}$ is nonempty if and only if $\mathfrak{a}(x^{lu})\geq\mathfrak{b}(x^{lu})$ (or equivalently $\theta_{\ell}(x^{lu})\geq-1$). Finally, for all $x\in X_{\ell}$ such that $\mathfrak{a}(x)\geq\mathfrak{b}(x)$, we have $\theta_{\ell}(x)\geq-1$. From  (21), we also have $|\lambda_{1}/\lambda_{2}|=\lambda_{1}/\lambda_{2}\geq\theta_{\ell}(x)$ for all $x\in X_{\ell}$ such that $\mathfrak{a}(x)\geq\mathfrak{b}(x)$.
As a corollary, we have
[…]
From the computation of $\overline{{\theta_{\ell}}_{X_{\ell}}}$, we define:
[…]
and
$\overline{\varepsilon}=\min\{\varepsilon^{<0},\varepsilon^{>0}\}.$ (30)
Finally, we have for all $x\in X_{\ell}$ and all $k\in\mathbb{N}$
$\varphi_{\ell}(L^{k}(x))\leq\gamma_{\overline{\varepsilon}}\left(\theta_{\ell}(x),k\right)\text{ where }\theta_{\ell}=-|\mathfrak{b}/\mathfrak{a}|.$ (31)
We insist on the fact that $\gamma_{\overline{\varepsilon}}\in\mathcal{KL}_{\rm gen}$ and $\theta_{\ell}\in\operatorname{Fin}\left(X_{\ell}\right)$. The details of the proof of (31) is provided in the Appendix.
From sufficiently large values of $t$, $\gamma_{\overline{\varepsilon}}(\overline{{\theta_{\ell}}_{X_{\ell}}},t)=H(\overline{{\theta_{\ell}}_{X_{\ell}}},|\lambda_{1}/\lambda_{2}|^{t})$ and thus $\inf_{t\geq 0}\gamma_{\overline{\varepsilon}}(\overline{{\theta_{\ell}}_{X_{\ell}}},t)=H\left(\overline{{\theta_{\ell}}_{X_{\ell}}},0\right)=\lambda_{2}/c$. It follows that $\{k\in{\mathrm{R}}_{\omega}^{s}:\inf_{t\geq 0}\gamma_{\overline{\varepsilon}}(\overline{{\theta_{\ell}}_{X_{\ell}}},t)<\omega_{k}\}={\mathrm{R}}_{\omega}^{s}$ and finally the pair $(\gamma_{\overline{\varepsilon}},\theta_{\ell})$ is useful.
Applying our computations to the numerical data of Subsection (3.4), we get
$\mathfrak{a}(x^{lu})=95>0\text{ and }\mathfrak{a}(x^{ul})=-10<0$
As $\mathfrak{b}(x^{lu})=55$ and $\mathfrak{b}(x^{ul})=25$, we have
$\overline{{\theta_{\ell}}_{X_{\ell}}}=\max\{\theta_{\ell}(x^{lu}),\theta_{\ell}(x^{ul})\}=\max\{-11/19,-2.5\}=-11/19$
and as $\mathfrak{a}(x^{lu})>\mathfrak{b}(x^{lu})$,
$\overline{\varepsilon}=\min\{0.5(-1+2.5),0.5(-0.5+11/19)\}=3/76$
Finally, we obtain:
$\varphi_{\ell}({L_{e}}^{k}x)\leq\gamma_{\overline{\varepsilon}_{e}}(\theta_{\ell}(x),0.5^{t})=\left\{\begin{array}[]{lr}4\dfrac{0.5^{k+1}+\theta_{\ell}(x)}{0.5^{k}+\theta_{\ell}(x)}&\text{ if }\theta_{\ell}(x)\leq-(0.5)^{t}-\dfrac{3}{76}\\ \\ \theta_{\ell}(x)+153\times 0.5^{t}+\dfrac{307}{76}&\text{ if }\theta_{\ell}(x)>-0.5^{t}-\dfrac{3}{76}\end{array}\right.$.
Following the notations introduced earlier, we have the following equivalence:
$\nu_{\rm opt}<+\infty\iff\sup_{k\in\mathbb{N}}\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right).$
We recall that:
$\nu_{\rm opt}<+\infty\iff\limsup_{k\to+\infty}\nu_{k}=\limsup_{k\to+\infty}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))=\limsup_{k\to+\infty}\ \overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}\in\mathbb{R}\cup\{-\infty\}$
We can establish the first result, which states that the existence of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds is equivalent to the fact that the sequence $\nu$ is bounded from above. This is an adaptation of classical results dealing with stability and $\mathcal{KL}$ upper bounds.
@@
The sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$ satisfies $\nu_{\rm opt}<+\infty$ if and only if there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. If, moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ then $(\gamma,\theta)$ can be chosen to be useful for $\nu$.